OpenStudy (lena772):

For the function f(x) = x + 7, what is the ordered pair for the point on the graph when x = 2b? (2b, 2b + 7) (2b, x + 7) (x, x + 7) (x, 2b + 7)

OpenStudy (lena772):

@jim_thompson5910 @Zale101

OpenStudy (zale101):

Replace all x with 2b f(x) = x + 7 f(2b) = 2b + 7

OpenStudy (zale101):

f(x) replaces y so (x,f(x))

OpenStudy (lena772):

So the first one?

OpenStudy (zale101):

f(2b) = 2b + 7 because x=2b we replace each x with 2b that leaves us with the the ordered pair (2b, 2b+7)

OpenStudy (lena772):

Thanks!

OpenStudy (zale101):

no problem, i'm really sorry about the previous question x_x

OpenStudy (lena772):

Don't worry about it. We're human. We're entitled to make mistakes sometimes. :)